MCV NL LAMIVI 20. Represent the following plane geometrically : 2x-y-5z = -2 (K/C)
Answer #1

Solh - given plane Tax-y-sz=-2. Step- ist in xey-plane yo. 160,2) 2=0 so 28-y=-2 at xao y=42 2x+y=- 27 (-1,0) 7 d atyzo x=-1